HIGH-RESOLUTION NEAR-INFRARED SPECTROSCOPY OF He/N_{2}/H_{2} POSITIVE-COLUMN PLASMAS

\underline{JENNIFER L. GOTTFRIED}, CHRISTOPHER P. MORONG, and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

Like its isoelectronic cousins BH_{2}^{−} and CH_{2}, the amidogen cation NH_{2}^{+} has a quasilinear ground state X̃^{3}B_{1} with a low barrier to linearity (155 cm^{−1}) and metastable excited electronic states ã^{1}A_{1} and b̃^{1}B_{1} that become degenerate (^{1}Δ) at linearity. In addition to its theoretical interest (due to the quasilinearity and the Renner effect), NH_{2}^{+} is one of the most fundamental molecular ions that exist abundantly in laboratory plasmas containing hydrogen and nitrogen. Despite this, only two high-resolution experimental detections of NH_{2}^{+} have been reported (the observation of the antisymmetric N-H stretch^{a} at 3360 cm^{−1}, and four hot bands^{b} from 2900-3500 cm^{−1}).

In an attempt to observe the predicted near-infrared electronic absorption spectrum of NH_{2}^{+},^{c} we have recently obtained new spectra of positive ions in a liquid-nitrogen-cooled positive column He/N_{2}/H_{2} plasma. The spectra were recorded using a high-resolution, high-sensitivity spectrometer based on a Ti:sapphire laser (11,000-13,000 cm^{−1}) and incorporating velocity modulation, phase modulation with heterodyne detection, noise subtraction, and optical multi-passing. The observation and assignment of the spectra is complicated by the presence of thousands of lines from the A^{2}Π_{u} − X^{2}Σ_{g}^{+} system of N_{2}^{+}. We will report the results of our analysis of the new spectra.

aM. Okumura, B. D. Rehfuss, B. M. Dinelli, M. G. Bawendi, and T. Oka, J. Chem. Phys. 90, 5918 (1989).

bY. Kabbadj, T. R. Huet, D. Uy, and T. Oka, J. Mol. Spectrosc. 175, 277 (1996).

cG. Osmann, P. R. Bunker, P. Jensen, and W. P. Kraemer, J. Mol. Spectrosc. 186, 319 (1997).